OpenStudy (anonymous):

On Monday morning, Colin filled his car with six and one-half gallons of gas, then used three and one-fourth gallons driving. On Tuesday, he used up 2 gallons driving. On Tuesday night, his gas tank had four and one-fourth gallons remaining. How much gas was in Colin's tank before he filled up on Monday morning? 3 gal three and one-fourth gal three and one-half gal three and three-fourths gal

OpenStudy (ganpat):

let x be previous fuel 6 1/2 gallons petrol filled.. 3 1/4 + 2 used fuel.. remaining gallons = 4 1/4.. so ,. x +13/4 = 2 + 9/4 + 13/2 multiply throughout by 4.. 4x + 13 = 8 + 9 + 26 4x = 30 x = 30/4 = 7 1/2 gallons... i guess...

OpenStudy (anonymous):

thx

OpenStudy (ganpat):

w/c.. :)

OpenStudy (anonymous):

No. The answer is 3 3/4 gallons. And there isn't a choich that says 7 1/2 gallons!